OpenStudy (anonymous):

Which of the following BEST describes how to use the addition property of equality to isolate the variable m below? –6 = –3 + m A. add 3 to both sides of the equation B. subtract 3 from both sides of the equation C. add 6 to both sides of the equation D. subtract 6 from both sides of the equation

OpenStudy (astrophysics):

Hey, guess I'll finish helping ya :)

OpenStudy (igreen):

You would add 3 to both sides..

OpenStudy (igreen):

Because: 3 is a negative, so you do the opposite. It will isolate m and give you m = -3

OpenStudy (astrophysics):

Ok so from what we've learned so far, we need to get m on it's own, to do this we need to add 3 on both sides, here I'll show you what happens when you do it. \[-6 +3 = -3+3 +m\] You see -3+3 = 0 so on the left side we're left with just m, and -6+3 is -3. \[m = 3\] So your answer is add by 3 :)